OpenStudy (adunb8):

linear algebra & differential equation question using bernoulli equation y'-2(sinx)y = -2y^3/2sinx

OpenStudy (lgbasallote):

it's already in the bernoulli form...so now..identify what y^n is

OpenStudy (anonymous):

how far along are you?

OpenStudy (adunb8):

um.. y^n would be y^3/2?

OpenStudy (anonymous):

I think it's supposed to be y^3 in your problem statement.

OpenStudy (adunb8):

? what do you mean \[y'-2(sinx)y = -2y ^{3/2}sinx\]

OpenStudy (adunb8):

i dont know what to do after

OpenStudy (anonymous):

ok just making sure..

OpenStudy (anonymous):

divide through by y^(3/2)

OpenStudy (anonymous):

make the sub.s y^(-1/2) =v

OpenStudy (anonymous):

you there?

OpenStudy (adunb8):

yes

OpenStudy (anonymous):

ok... now find v' , remember to use the chain rule...

OpenStudy (adunb8):

wait where did you get y^1/2?

OpenStudy (anonymous):

v= y^(1-n) n is 3/2

OpenStudy (anonymous):

is that what you're asking about?

OpenStudy (adunb8):

yes oh isee you divide all of them with y^3/2

OpenStudy (anonymous):

cool. So what did you get for v'?

OpenStudy (adunb8):

i got y^-3/2 - 2(sinx)y^-v = -2sinx

OpenStudy (adunb8):

y' on the y^3/2

OpenStudy (anonymous):

v' = -1/2 * y^(-3/2) * y'

OpenStudy (anonymous):

now make the sub.s so all the y and y' are in terms of v and v'

OpenStudy (adunb8):

\[y ^{-3/2}y' -2(sinx)y ^{-1/2} = -2sinx\]

OpenStudy (adunb8):

this is my step right now

OpenStudy (anonymous):

y^(-1/2) = v y^(-3/2) * y' = -2*v'

OpenStudy (anonymous):

now you have an equation in v, v' and x ... you can solve for v using the method of 'integrating factor' or whatever else you are comfortable with.

OpenStudy (adunb8):

so you changed the equation into this? what happened to the sinx?

OpenStudy (anonymous):

nothing

OpenStudy (adunb8):

can you show me the full equation what you changed im not quite getting it..

OpenStudy (anonymous):

show me what you have...

OpenStudy (adunb8):

\[y ^{-3/2}y' -2(sinx)y ^{-1/2} = -2sinx\]

OpenStudy (anonymous):

y^(-1/2) = v y^(-3/2) * y' = -2*v'

OpenStudy (adunb8):

so it would be -2v' - 2(sinx)v = -2sinx ??

OpenStudy (anonymous):

v' +(sinx)v = sinx

OpenStudy (adunb8):

oh isee then use the 1st order to solve? or homogenous?

OpenStudy (anonymous):

"backspace" is sticking lol

OpenStudy (anonymous):

yep:)

OpenStudy (adunb8):

that is confusing about bernoulli is am i allow to just change y^(-3/2)y' into just v' ? and how did we get the 2v' i get the v' part but how did we get the 2?

OpenStudy (adunb8):

how did you know to make it y^(-3/2) * y' = -2*v'

OpenStudy (adunb8):

-2? from

OpenStudy (anonymous):

y^(-1/2) =v -1/2 * y^(-3/2) * y' = v'

OpenStudy (anonymous):

just took the derivative of the "v" we chose...

OpenStudy (anonymous):

all good?

OpenStudy (adunb8):

yes i think i got it we took the derivative because its v' right?

OpenStudy (anonymous):

we took the derivative because we're trying to substitute in a simpler expression for y' and any (and all) powers of y. Can't do that unless we have v and v'.

OpenStudy (adunb8):

one last question can it be different variable? or it has to be v and v'?

OpenStudy (anonymous):

It can be whatever you like as long as it's not x :)

OpenStudy (adunb8):

thank you~~~~~

OpenStudy (anonymous):

Sure:)